# PhysGM: Large Physical Gaussian Model for Feed-Forward 4D Synthesis

### v1

@@ Conclusion
Automated Ranking and Preference Pair Construction
To automate the preference labeling process, we implement a quantitative comparison method. We employ CoTracker-3, a state-of-the-art model for dense point tracking, to trace the trajectories of a predefined set of pixels across the ground-truth video and each of the three candidate videos. The physical and perceptual similarity between each candidate and the ground truth is quantified by calculating the L2 distance (Euclidean distance) between their corresponding pixel trajectories. A lower L2 distance signifies a closer match to the ground-truth motion, indicating higher physical fidelity. Based on the calculated L2 distances, we establish a definitive preference ranking (e.g., Candidate A >Candidate B >Candidate C, where A has the lowest L2 distance). This ranked list is subsequently decomposed into multiple (winner, loser) preference pairs. The resulting collection of pairs forms the final dataset that directly fuels our DPO fine-tuning, enabling the model to discern and internalize the subtle principles governing realistic physical dynamics.
Additional Data Sources
@@
It is also worth noting that the PhysX (Cao et al. 2025) dataset was released concurrently with our research, offering 3D objects annotated with physical properties ,which is also suitable for our dataset process. Given the timing constraints, its integration was not feasible for the present study. Nevertheless, we acknowledge its significance and view it as a promising avenue for extending our work in the future.
@@
Additional Results
To demonstrate the versatility and robustness of our approach, we present a range of additional experiments. We showcase qualitative results for fundamental stretching and dropping scenarios in Figure 6 and Figure 7, respectively. These experiments are further extended to demonstrate our model’s performance across more challenging configurations: Figure 8 illustrates its robustness in scenes with complex backgrounds, while Figure 9 highlights its capability to handle multi-object interactions.Additional results are presented in the Figure 10 Figure 11 Figure 12.
Limitations and Future Work
Data Dependency and Generalization.
Simplified Physics Representation.
Limitations of Base Models.
Our method for 3D reconstruction primarily relies on LGM and MVDream as its foundational models. While these models perform effectively in most scenarios, their inherent limitations can still impact the quality of the final output, potentially leading to issues such as the loss of geometric detail, the introduction of artifacts, or textural inconsistencies. Consequently, a key direction for future work is to explore and integrate more advanced and robust generative models as replacements.

### v2

@@ D.2.3 Candidate Video Generation
D.2.4 Preference Labeling via Trajectory Alignment
To create preference pairs for DPO training, we develop an automatic labeling pipeline that compares simulated dynamics against reference videos through three-stage trajectory alignment. We employ SAM-2 [ravi2024sam] for object segmentation and CoTracker-3 [karaev24cotracker3] for dense trajectory extraction across both ground-truth and simulated sequences.
Spatial Alignment.
Due to different camera viewpoints and object scales between reference and simulated videos, direct trajectory comparison is infeasible. We address this through bounding box normalization: for each video, we compute the object’s bounding box from its segmentation mask as $\mathcal{B}=(x_{\min},y_{\min},x_{\max},y_{\max})$. Point trajectories from the ground-truth video are first normalized to $[0,1]$ coordinates relative to its bounding box:
$\mathbf{p}_{\text{norm}}=\left(\frac{x-x_{\min}}{x_{\max}-x_{\min}},\frac{y-y_{\min}}{y_{\max}-y_{\min}}\right)$ (15)
These normalized coordinates are then mapped to the simulated video’s coordinate frame using its bounding box parameters. This spatial alignment ensures correspondence between trajectories regardless of viewpoint or scale differences.
Landing Frame Alignment.
Physical simulations may exhibit different temporal dynamics (e.g., falling speeds) even with similar physical properties. To enable fair comparison, we align sequences based on a key physical event: the object’s landing moment. Specifically, we identify the landing frame as the temporal turning point where the object’s vertical motion reverses. For each video, we track the point with maximum $y$-coordinate in the first frame (typically the object’s bottom) and monitor its trajectory. The landing frame $f^{*}$ is detected when the vertical velocity changes sign:
$f^{*}=\arg\min_{t}\{t\mid y_{t}\leq y_{t-1},t>0\}$ (16)
where $y_{t}$ represents the tracked point’s $y$-coordinate at frame $t$. This frame marks the transition from falling to resting/bouncing phases.
Temporal Alignment.
Using the detected landing frames $(f^{*}_{\text{GT}},f^{*}_{\text{sim}})$ as temporal anchors, we align the post-landing phases of both sequences. We determine the comparable duration as $T=\min(T_{\text{GT}}-f^{*}_{\text{GT}},T_{\text{sim}}-f^{*}_{\text{sim}})$, where $T_{\text{GT}}$ and $T_{\text{sim}}$ are the total frame counts. Additionally, we compute a spatial offset $(\Delta x,\Delta y)$ between the landing positions in both videos and apply this correction to the simulated trajectories:
$\mathbf{p}_{\text{sim}}^{\text{aligned}}=\mathbf{p}_{\text{sim}}+(\Delta x,\Delta y)$ (17)
This ensures that both sequences are aligned not only temporally but also spatially at the critical landing event.
Similarity Metric.
After three-stage alignment, we compute the trajectory dissimilarity as:
$d(\mathcal{V}_{\text{sim}},\mathcal{V}_{\text{GT}})=\frac{1}{NT}\sum_{n=1}^{N}\sum_{t=1}^{T}\|\mathbf{p}_{n,t}^{\text{GT}}-\mathbf{p}_{n,t}^{\text{sim}}\|_{2}$ (18)
where $N$ is the number of tracked points and $T$ is the aligned sequence length. Lower dissimilarity indicates better physical plausibility. For each scene, we rank $K$ candidate simulations by this metric and select the best as the “winner” and worst as the “loser” for DPO training.
D.2.5 Additional Data Sources
@@
It is also worth noting that the PhysX [cao2025physx] dataset was released concurrently with our research, offering 3D objects annotated with physical properties ,which is also suitable for our dataset process. Given the timing constraints, its integration was not feasible for the present study. Nevertheless, we acknowledge its significance and view it as a promising avenue for extending our work in the future.
Figure 8: Multi-view generation using MVAdapter. Given a single frontal view image as input (left), MVAdapter [huang2025mv] generates three auxiliary views: rear, left, and right (right three panels). These synthesized views, together with the input frontal view, provide comprehensive angular coverage for our 3D Gaussian reconstruction and physics prediction pipeline. The generated views maintain consistent geometry and appearance while capturing different perspectives of the object.
Appendix E Additional Results
To fully demonstrate the versatility and effectiveness of our approach, we present an extended suite of supplementary experiments with comprehensive qualitative insights. Specifically, Figure 9 provides detailed visualizations of the Multi-View Stereo (MVS) module, showcasing its exceptional ability to accurately reconstruct 3D geometry from multi-view inputs. Given four randomly selected input views, the model generates novel viewpoints, and we visualize four representative views sampled from eight randomly selected output views as examples. Complementarily, Figure 8 offers visualizations of the MVAdapter component, clearly revealing how it effectively bridges domain gaps and enhances feature alignment across diverse input modalities. Beyond the core component validations, Figure 10 and Figure 11 exhibit the model’s performance on fundamental stretching and dropping scenarios, respectively. We further push the envelope to validate its effectiveness under more challenging configurations: Figure 12 illustrates strong robustness in cluttered/complex background scenes, while Figure 13 highlights its superior capability in handling intricate multi-object interactions and other results in Figure 14 15 16.
Appendix F Limitations and Future Work
Data Dependency and Generalization.
Simplified Physics Representation.
Figure 9: Qualitative results for Multi-View Stereo. Our method generates Gaussian splatting with remarkable visual quality on various challenging images.